OpenStudy (anonymous):

What is 19 and 7/8% as a fraction or mixed number in its simplest form?

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

19+(7/8) = ....

OpenStudy (skullpatrol):

@HistoryGirl420 Are you still there?

OpenStudy (anonymous):

Yes I am @skullpatrol

OpenStudy (skullpatrol):

@HistoryGirl420 You should ask questions

OpenStudy (anonymous):

Kind of odd trick, but it works for this fraction. ((19*8)+7)/8 = 159/8

OpenStudy (anonymous):

Won't work for most fraction

OpenStudy (anonymous):

Explain how you got that answer in steps please? @larryboxaplenty

OpenStudy (anonymous):

There is a better way of doing it

OpenStudy (skullpatrol):

What is the definition of percent @HistoryGirl420 ?

OpenStudy (anonymous):

Percent means per hundred

OpenStudy (anonymous):

@skullpatrol

OpenStudy (skullpatrol):

Correct, so divide by 100.

OpenStudy (anonymous):

How do I divide with a mixed number?

OpenStudy (skullpatrol):

Treat it like any other single number. Just put a fraction bar under it and 100 in the denominator.

OpenStudy (anonymous):

Huh?

OpenStudy (skullpatrol):

$$\Huge 19\frac{7}{8}\text{%}$$

OpenStudy (skullpatrol):

$$\Huge \text{%} = \frac{1}{100}$$

OpenStudy (skullpatrol):

Ok so far?

OpenStudy (skullpatrol):

any questions?

OpenStudy (skullpatrol):

$$\Huge (19\frac{7}{8})\text{%}$$

OpenStudy (skullpatrol):

$$\Huge \frac{19\frac{7}{8}}{100}$$

OpenStudy (skullpatrol):

Treat it like any other single number. Just put a fraction bar under it and 100 in the denominator.

OpenStudy (skullpatrol):

@HistoryGirl420 Are you still there?

OpenStudy (anonymous):

Sorry. Had to take my dog out...